# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v1

@@ 4.2 Coder-Evaluator Framework
where denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases. Given the inherent hallucination tendencies of LLMs, candidate functions might contain logical or dimensionality errors. To address this, each candidate TFCAF undergoes a validation step by testing it with state information and local Q-values from a representative environment scenario. If an error occurs (such as dimension mismatch or logical inconsistencies), the error message is fed back to , prompting regeneration of the faulty TFCAF:
This correction step repeats until all candidate TFCAFs pass basic execution checks without errors.
@@
Subsequently, evaluates these error-free candidate TFCAFs, selecting the most promising one, while also providing specific programming tips and insights for further optimization:
where . The selection and programming tips produced by are incorporated into a new prompt for , guiding it to generate another candidate functions:
where . Through repeated iterations of this coder-evaluator interaction, the candidate functions are progressively refined until convergence to an optimized final TFCAF . This final function explicitly includes a state-dependent weights module and a bias module .
Iterative Refinement.
Ultimately, is integrated into a predefined TFCAF template (see Appendix A), replacing the traditional mixing network and effectively performing credit assignment across multiple agents in MARL scenarios.
4.3 Training Procedure
@@
5 Experiments
To comprehensively evaluate the proposed method, the main experiments were designed with four main objectives: (1) to assess the performance of the QLLM algorithm in terms of average episode return and variance across multiple environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other value decomposition algorithms; and (4) to validate the effectiveness of the proposed coder-evaluator framework under various conditions.
5.1 Experimental Setups
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary large language model used in our method is the DeepSeek-R1 [7] inference model. For each experiment, the reported return values represent the mean over five independent training runs, with 95% confidence intervals indicated by error bars. Unless otherwise specified, all environment parameters are set to their default values. Our experimental test environments include: (1) Cooperative Matrix Game111https://github.com/uoe-agents/matrix-games: This is a non-monotonic cooperative case where two cooperating agents must find an optimal Nash equilibrium [44]. (2) Multi-agent Particle Environments (MPE)222https://pettingzoo.farama.org/environments/mpe/: This environment consists of a series of cooperative tasks involving multiple particles, including navigation, predation, and communication. Agents must also avoid penalties caused by collisions. (3) Level-Based Foraging (LBF)333https://github.com/uoe-agents/lb-foraging: Multiple agents must collaborate to collect all the food items in the map. Successful collection occurs only when the sum of the levels of the participating agents is greater than or equal to the level of the food. (4) Robotic Warehouse (RWARE)444https://github.com/uoe-agents/robotic-warehouse: Robots continuously receive instructions to retrieve goods from specified locations and transport them to target locations. The detailed parameter settings, reward settings, and global state-space settings for each environment can be found in Appendix B.
We selected five commonly used baseline algorithms for comparison in this study:
Value Decomposition Algorithms: QMIX [28] is a value-based MARL algorithm that learns a joint Q-value function by factorizing individual agent value functions while maintaining a monotonic relationship between them. This allows agents to work independently, yet still guarantee optimal joint actions.
Policy Gradient Algorithms COMA [6] is a policy-gradient-based approach that uses counterfactual reasoning to address the credit assignment problem in multi-agent settings. It aims to improve the learning of individual agents by considering the effect of each agent’s action on the overall outcome. MADDPG [19] extends the DDPG algorithm to multi-agent settings. It utilizes centralized training with decentralized execution, where each agent learns a policy conditioned on the global state during training, but acts based on its local state during execution.
Reward Interpolation-Based Algorithms: RIIT [10] is a reward-based algorithm that improves multi-agent learning by interpolating rewards across agents. This helps agents better coordinate their actions to achieve a common objective.
Attention-Based Algorithms: Qatten [43] integrates attention mechanisms into Q-learning, allowing agents to dynamically adjust the focus on different aspects of the environment. This improves coordination and allows the agents to make more informed decisions.
5.2 The Superiority of QLLM
@@
Figure 4 and Figure 5 shows the performance of various algorithms in certain scenarios. It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward LBF environment. Especially in the LBF environment, COMA, MADDPG, and RIIT almost fail to train the agents properly, which highlights the importance of accurate credit assignment in sparse-reward environments. In Listing 1, we present the credit assignment function generated by the large language model for the 2s-10x10-3p-3f map in the LBF environment. It is evident that this function takes into account key factors such as the positions of the agents and food, the effectiveness of the food, and the actions of the agents, accurately assigning contributions to each agent. Furthermore, it is easily extendable to similar scenarios.
@@
We also tested the performance of various algorithms in high-dimensional state spaces in certain scenarios. Taking the simple-spread task in MPE environment as an example, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially. From the Figure 6, it can be seen that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in the QLLM. The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult. However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
@@
Our algorithm is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values. We replaced the mixing network in RIIT with the TFCAF generated by QLLM and tested the impact of using TFCAF on the performance of RIIT in three environment of MPE. The experimental results are shown in Table 1. Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
Environment | RIIT-QLLM | RIIT |
pz-mpe-simple-spread-N6 | -169.7 3.4 | -201.1 2.7 |
pz-mpe-simple-adversary | 101.5 1.8 | 96.3 0.9 |
pz-mpe-simple-tag | 217.2 1.3 | 195.7 0.7 |
5.3 Ablation Studies
5.3.1 Coder-Evaluator Framework’s Candidate Responses
To study the impact of the number of candidate responses, a key hyperparameter in QLLM, Figure 7 shows the post-convergence average episode return of QLLM with varying on the tasks simple-tag and lbf:2s-10x10-3p-3f. The results indicate that as long as a sufficient number of candidate responses are provided, QLLM is insensitive to this hyperparameter. This finding supports our claim that the coder-evaluator framework helps reduce the randomness in LLM inference. Therefore, we set for all tasks in this work.
5.3.2 Evaluator LLM’s Effectiveness
@@
Most RL algorithms enhanced by large language models use only one kind of LLM to do his job directly. Our coder-evaluator framework features two large language models, each with its specific role, working together to generate an accurate credit assignment function. Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in large language models. We test the performance of QLLM across different reinforcement learning environments without using the evaluator LLM. Detailed experimental details can be found in Appendix B.
6 Conclusion
In this study, we propose QLLM, a method that leverages large language models to address the credit assignment problem in multi-agent reinforcement learning. Our algorithm employs a coder-evaluator framework to generate training-free, task-specific credit assignment functions. This frame significantly reduces hallucination and laziness in code generation by large language models, while also improving the accuracy of task understanding.
Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, the proposed algorithm exhibits strong generalization capability and is compatible with various MARL algorithms that employ mixing networks, making it a promising and generalizable solution for complex multi-agent scenarios.
Limitations & Future Works: A current limitation of this work lies in the need to manually craft prompts for reinforcement learning tasks, which can become challenging due to the complexity of certain tasks. As LLM capabilities continue to advance, prompt design in QLLM is expected to become easier, requiring less task-specific information and further reducing manual effort. Future work will focus on expanding the applicability of the coder-evaluator framework and deploying the QLLM algorithm in real-world multi-robot systems.

### v4

@@ A Training-Free Credit Assignment Function (TFCAF) is defined as:
where denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases.
@@
Due to the inherent hallucination tendencies of LLMs, candidate functions might contain syntax or dimensional errors. To address this, each candidate TFCAF is compiled and run once with any available input state and local Q-values. If an error occurs during execution, the error message generated by the compiler will be fed back to , prompting regeneration of the faulty TFCAF:
This correction step repeats until all candidate TFCAFs can execute successfully without errors.
Selection of TFCAF.
Subsequently, evaluates these validated candidate TFCAFs, selecting the most promising one based on task relevance, interpretability, and code clarity. Notably, the selection of the best candidate is fully guided by structured prompts and performed automatically by , without any manual or subjective intervention:
where .
Training Procedure
@@
Experiments
To comprehensively evaluate the proposed method, the experiments were designed with six main objectives: (1) to assess the performance of the QLLM algorithm in terms of average return across multi-agent environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other mixing-based algorithms; (4) to demonstrate the compatibility of QLLM with various LLMs; (5) to validate the effectiveness of the proposed coder-evaluator framework under various parameters; (6) to demonstrate the superiority of our method in terms of training costs.
Experimental Setups
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary LLM used in our method is the DeepSeek-R1 (Guo et al. 2025) inference model. For each experiment, the reported return values represent the mean over five independent training runs, with 90% confidence intervals indicated by error bars. Unless otherwise specified, all environment parameters are set to their default values. Our experimental environments include: (1) Cooperative Matrix Games111https://github.com/uoe-agents/matrix-games, used to validate the algorithm’s performance in non-monotonic scenarios, which may violate the IGM principle; (2) Multi-agent Particle Environments (MPE)222https://pettingzoo.farama.org/environments/mpe/, for evaluating performance in high-dimensional state spaces; (3) Level-Based Foraging (LBF)333https://github.com/uoe-agents/lb-foraging, for testing in sparse reward settings; and (4) Google Research Football (GRF)444https://github.com/google-research/football, to assess performance in complex, realistic environments.
We compare QLLM with six commonly used baseline algorithms: QMIX (Rashid et al. 2020b), QPLEX (Wang et al. 2020), Weighted QMIX (Rashid et al. 2020a), Qatten (Yang et al. 2020), RIIT (Hu et al. 2021), MASER (Jeon et al. 2022), and COMA (Foerster et al. 2018). A detailed introduction and implementation details for all baselines and environments are provided in Appendix B.
The Superiority of QLLM
@@
Figure 3 and Figure 4 show the performance of various algorithms in certain scenarios. It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward GRF environment. Especially in the GRF environment, COMA almost fail to train the agents properly, which highlights the importance of accurate credit assignment in multi-agent environments. Furthermore, the incorporation of the IGM-Gating mechanism enables our algorithm to perform effectively in non-monotonic matrix game environments, outperforming methods like QMIX that are limited by rigid monotonicity constraints. In Listing 1, we present an example of a credit assignment function generated by LLMs for the pz-mpe-simple-spread task in the MPE environment. It can be seen that this function takes into account key factors such as the agents’ positions, their distances to landmarks, as well as the number of collisions, accurately assigning contributions to each agent. These verify the effectiveness and generalization ability of the credit assignment functions generated by our method.
@@
We tested the performance of various algorithms in high-dimensional state spaces of certain scenarios. Taking the pz-mpe-simple-spread task as an example in the MPE environment, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially. In Figure 6, the results indicate that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in QLLM. The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult. However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
@@
QLLM is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values. We used the TFCAF generated by QLLM to replace the mixing networks in three value-decomposition algorithms—RIIT, MASER, and QMIX. We then conducted experiments in the MPE environment to examine how integrating TFCAF affects the performance of these value-decomposition MARL methods. The experimental results tested in MPE environments are shown in Figure 5. Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
Conclusion
In this study, we propose QLLM, a method that uses LLMs to address the credit assignment problem in multi-agent reinforcement learning through coder-evaluator framework. This approach generates training-free, task-specific credit assignment functions, reducing hallucination and improving task understanding. Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-based algorithms. Future work will explore broader applications and deployment in real-world multi-robot systems.
Appendix A LLM Prompts and Responses
Below are the prompt template, example task information, and LLM response in our work.
Appendix B Details of Environments and Baselines
Baseline Algorithms
QMIX (Rashid et al. 2020b) is a representative value decomposition method designed for cooperative MARL. It factorizes the joint action-value function into individual agent utilities through a mixing network that enforces a monotonic constraint, ensuring that maximizing local utilities also maximizes the global return. The mixing network parameters are generated by a hypernetwork conditioned on the global state, enabling state-dependent flexible mixing.
QPLEX (Wang et al. 2020) builds on QMIX by introducing a duplex dueling architecture to overcome the limited expressiveness caused by strict monotonicity. It splits the value factorization into a main (advantage) stream and a complementary stream, allowing more flexible interactions while maintaining decentralized policies. This design expands the class of joint value functions that can be represented under the CTDE paradigm.
Weighted QMIX (Rashid et al. 2020a) further generalizes QMIX by learning an additional weighting factor for each agent’s contribution to the global Q-value. By relaxing the monotonicity assumption through these learnable weights, it improves the accuracy of credit assignment in tasks with more complex cooperative dependencies.
Qatten (Yang et al. 2020) enhances value decomposition by embedding a multi-head attention mechanism into the mixing network. This mechanism adaptively adjusts the importance of each agent’s local utility based on the global state context, enabling the model to focus on the most relevant agents or features when computing the global Q-value.
RIIT (Hu et al. 2021) tackles the multi-agent credit assignment problem by modeling implicit influence relationships among agents. It proposes a reward interpolation technique that redistributes the global team reward across agents, considering both immediate and long-term interdependencies, which helps reduce misattribution and improves coordination performance.
COMA (Foerster et al. 2018) is a centralized actor-critic method designed for cooperative MARL, with a focus on addressing the multi-agent credit assignment problem. It employs a centralized critic that estimates a counterfactual baseline for each agent by marginalizing out its action while keeping others fixed, enabling the computation of an advantage function that isolates each agent’s contribution. This counterfactual reasoning helps reduce variance in the policy gradient and promotes coordinated behaviors.
MASER (Jeon et al. 2022) is a value-decomposition-based MARL algorithm designed to improve credit assignment under sparse and delayed rewards. It introduces subgoals automatically generated from the experience replay buffer, which are used to construct intrinsic rewards that guide agents toward meaningful intermediate states. By augmenting standard Q-learning with subgoal-conditioned value estimation, MASER enhances exploration efficiency and coordination while remaining compatible with centralized training and decentralized execution.
Code Repository
The baseline algorithm in this experiment is implemented based on the epymarl555https://github.com/uoe-agents/epymarl and pymarl2666https://github.com/hijkzzz/pymarl2 codebases. We tested MPE, LBF, and matrix games using the epymarl library, while Google Research Football was evaluated with pymarl2. Although the SMAC777https://github.com/oxwhirl/smac environment is a popular benchmark for multi-agent reinforcement learning and is built upon PySC2888https://github.com/google-deepmind/pysc2, it lacks comprehensive documentation regarding the specific unit types, abilities, and map configurations used in each scenario. This obscurity hinders reproducibility and limits fine-grained control over experimental variables. Consequently, we opted for several environments with better transparency and configurability to ensure experimental rigor and facilitate controlled evaluations.
Tasks Details
Matrix Games
The cooperative matrix games environment is a simple benchmark setting in multi-agent reinforcement learning, where two or more agents simultaneously select actions from predefined action spaces. The environment is defined by a payoff matrix that specifies the global reward received based on the joint actions of all agents (below is the payoff matrix used in test environment). Cooperation is essential, as the optimal strategy often depends on coordinated decision-making among agents. In order to make sense of the global state space of this task, we include the actions executed by the agents last step.
Level-Based Foraging
The Level-Based Foraging environment consists of tasks focusing on the coordination of involved agents. The task for each agent is to navigate the grid-world map and collect items. Each agent and item is assigned a level and items are randomly scattered in the environment. In order to collect an item, agents have to choose a certain action next to the item. However, such collection is only successful if the sum of involved agents’ levels is equal or greater than the item level. Agents receive reward equal to the level of the collected item. The tasks that were selected are denoted first by the grid-world size (e.g. 15 × 15 means a 15 by 15 grid-world). Then, the number of agents is shown (e.g. “4p” means four agents/players), and the number of items scattered in the grid (e.g. “3f” means three food items). We also have some special flags, the “2s” which denotes partial observability with a range of two squares. In those tasks, agents can only observe items or other agents as long as they are located in a square of size 5x5 centred around them. Finally, the flag “c” means a cooperative-only variant were all items can only be picked up if all agents in the level attempt to load it simultaneously. Since the global state space of the original environment is stitched together from the observation vectors of individual agents, it contains a lot of duplicated information and some environmental information is lost. Therefore, we redesigned a global state space which contains all the environment information while reducing the size of the dimension. Specifically, it includes the coordinates and levels of individual intelligences and things, as well as the actions executed by each agent last step.
| LBF tasks | n_agents | state shape | action num |
| Foraging-2s-15x15-4p-3f | 4 | 25 | 6 |
| Foraging-2s-8x8-2p-2f-coop | 2 | 14 | 6 |
| Foraging-2s-10x10-3p-3f | 3 | 21 | 6 |
Multi-Agent Particle Environments
Simple-Spread: In this task, three agents are trained to move to three landmarks while avoiding collisions with each other. All agents receive their velocity, position, relative position to all other agents and landmarks. The action space of each agent contains five discrete movement actions. Agents are rewarded with the sum of negative minimum distances from each landmark to any agent and a additional term is added to punish collisions among agents.
Simple-Adversary: In this task, two cooperating agents compete with a third adversary agent. There are two landmarks out of which one is randomly selected to be the goal landmark. Cooperative agents receive their relative position to the goal as well as relative position to all other agents and landmarks as observations. However, the adversary agent observes all relative positions without receiving information about the goal landmark. All agents have five discrete movement actions. Agents are rewarded with the negative minimum distance to the goal while the cooperative agents are additionally rewarded for the distance of the adversary agent to the goal landmark. Therefore, the cooperative agents have to move to both landmarks to avoid the adversary from identifying which landmark is the goal and reaching it as well. For this competitive scenario, we use a fully cooperative version where the adversary agent is controlled by a pretrained model.
Simple-Tag: In this task, three cooperating predators hunt a forth agent controlling a faster prey. Two landmarks are placed in the environment as obstacles. All agents receive their own velocity and position as well as relative positions to all other landmarks and agents as observations. Predator agents also observe the velocity of the prey. All agents choose among five movement actions. The agent controlling the prey is punished for any collisions with predators as well as for leaving the observable environment area (to prevent it from simply running away without needing to learn to evade). Predator agents are collectively rewarded for collisions with the prey. We employ a fully cooperative version of this task with a pretrained prey agent.
| MPE tasks | n_agents | state shape | action num |
| pz-mpe-simple-spread | 6 | 216 | 5 |
| pz-mpe-simple-tag | 3 | 48 | 5 |
| pz-mpe-simple-adversary | 2 | 16 | 5 |
Google Research Football
The Google Research Football environment is a challenging reinforcement learning platform designed to simulate football matches in a highly realistic and strategic setting. Agents control individual players and must learn to cooperate, plan, and execute complex behaviors to achieve objectives such as scoring goals or defending against opponents. The environment provides various predefined scenarios as well as full 11-versus-11 matches, supporting both single-agent and multi-agent training. In the Google Research Football environment, the action space consists of a discrete set of predefined high-level actions such as moving, passing, shooting, and sprinting, allowing agents to focus on strategic decision-making rather than low-level motor control. The Simple115 observation space is a fixed-size vector with 115 features, providing essential information about the game state, including the positions and velocities of the ball and players, the directions of movement, and other key attributes necessary for learning effective policies.
| GRF tasks | n_agents | state shape | action num |
| academy_3_vs_1_with_keeper | 4 | 460 | 19 |
| academy_counterattack_easy | 4 | 460 | 19 |
| academy_pass_and_shoot_with_keeper | 3 | 345 | 19 |
Appendix C Additional Experiments
Compatibility Testing of Different LLMs
We tested the performance of QLLM with three mainstream LLMs, Deepseek-R1, Deepseek-V3999https://www.deepseek.com/, and ChatGPT-4o101010https://openai.com/, along with other algorithms in the MPE environment. As shown in Figure 8, the experimental results indicate that mainstream LLMs outperform the baseline algorithms. Moreover, as the inference capabilities of the LLMs improve, the performance of our algorithm also improves.
Tests of the Effect of the Evaluator LLM
@@ Appendix C Additional Experiments
Most reinforcement learning algorithms enhanced by LLMs use only one kind of LLM to do their job directly. Our coder-evaluator framework features two LLMs, each with its specific role, working together to generate an accurate credit assignment function. Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in LLMs. We tested the performance of QLLM across different reinforcement learning environments without using the evaluator LLM. The experimental results (Figure 9) show that, the evaluator LLM significantly improves the algorithm’s performance in most environments.
Comparison of Algorithms Training Costs
As mentioned in the main text, QLLM replaces the mixing network with a deterministic functional expression, TFCAF, resulting in a significantly smaller number of learnable parameters. Therefore, the training cost of our algorithm is reduced. Table 4 lists the number of trainable parameters for each algorithm in the environments tested in this paper. Our algorithm reduces the number of trainable parameters by approximately 12% to 37%. This reduction directly lowers the training cost and simplifies the model structure, making the algorithm more efficient and improving performance.
| Environment / Algorithm | QLLM | QMIX | QATTEN | MADDPG | RIIT | QPLEX | Baseline Average | Reduction Rate (%) |
| pz-mpe-simple-spread | 105.221 | 140.422 | 191.114 | 154.246 | 191.175 | 154.762 | 166.744 | 36.88% |
| pz-mpe-simple-tag | 102.277 | 112.806 | 129.034 | 127.494 | 132.199 | 128.010 | 125.909 | 18.78% |
| pz-mpe-simple-adversary | 101.381 | 107.270 | 118.282 | 122.246 | 124.359 | 122.762 | 118.584 | 14.52% |
| Foraging-2s-15x15-4p-3f | 103.174 | 111.815 | 121.963 | 126.727 | 130.536 | 127.372 | 123.283 | 16.30% |
| Foraging-2s-8x8-2p-2f-coop | 101.766 | 106.887 | 116.683 | 122.119 | 124.296 | 122.764 | 118.150 | 13.87% |
| Foraging-2s-10x10-3p-3f | 102.662 | 109.735 | 120.043 | 124.807 | 127.912 | 125.452 | 121.190 | 15.28% |
| climbing-nostate-v0 | 100.099 | 103.684 | 110.408 | 118.148 | 118.661 | 118.406 | 113.861 | 12.07% |
| academy_3_vs_1_with_keeper | 430.355 | 529.300 | 601.496 | 634.900 | 568.085 | 543.014 | 575.359 | 25.23% |
| academy_counterattack_easy | 430.355 | 529.300 | 601.496 | 634.900 | 568.085 | 543.014 | 575.359 | 25.23% |
| academy_pass_and_shoot_with_keeper | 430.099 | 504.884 | 560.760 | 600.084 | 530.869 | 523.046 | 543.129 | 20.83% |